# Invariants of persistence modules defined by order-embeddings

### v2

@@ 5.1. Interval modules and the intermediate extension functor
    Consider $\mathcal{X}=\mathcal{X}^{\prime}_{3}$ or $\mathcal{X}^{\prime\prime}_{3}$ as follows
[…]
    We recall the description of the indecomposable $\mathcal{X}_{3}^{\prime}$- and $\mathcal{X}_{3}^{\prime\prime}$-modules from the proof of Proposition 3.12and Example 4.10: the modules $\Theta_{f}(U)$ are either simples, rectangles, or interval modules of the form ${\mathbb{I}}_{[f(1),f(2)]\cup[f(1),f(3)]}$ when $\mathcal{X}=\mathcal{X}^{\prime}_{3}$, and ${\mathbb{I}}_{[f(1),f(3)]\cup[f(2),f(3)]}$ when $\mathcal{X}=\mathcal{X}^{\prime\prime}_{3}$:
@@ 5.2. Families of embeddings
Let $\mathcal{P}$ be a poset. A family of $\mathcal{P}$-order embeddings is a finite set :=$\mathcal{F}:=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ with each $\mathcal{X}_{i}$ a finite poset (not necessarily representation finite) and each $\mathrm{E}_{i}$ a subset of $\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}}$. For =$\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ a family of $\mathcal{P}$-order embeddings, we denote by $\mathrm{E}:=\bigsqcup_{i}\mathrm{E}_{i}$ and we define an invariant $\mathbf{mult}_{\mathcal{F}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}}$ by
$\left(\mathbf{mult}_{\mathcal{F}}(M)\right)_{f}:={\rm mult}(\mathbb{I}_{\mathcal{X}_{i}},f^{*}(M))\quad\textrm{for }f\in\mathrm{E}_{i}.$
@@
    Let :=$\mathcal{F}:=\{(\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}})\}$ for $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as in Example 3.7. Then $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}\sqcup\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ is in natural bijection with $\{\leq_{\mathcal{P}}\}$ and via this bijection we have $\mathbf{mult}_{\mathcal{F}}=\operatorname{\mathbf{rk}}_{\mathcal{P}}$.
@@
      Let $\mathfrak{I}$ be the set of posets which embed into $\mathcal{P}$ as intervals, and for each $\mathcal{X}_{i}\in\mathfrak{I}$ let $\mathrm{E}_{i}\subseteq\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}}$ be the set of embeddings with interval image. Then, as a result of [CL18, Lemma 3.1], the generalized rank invariant of [KM21] is equal to $\mathbf{mult}_{\mathcal{F}}$ for =$\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\mid\mathcal{X}_{i}\in\mathfrak{I}\}$.
@@
      Suppose that $\mathcal{P}$ is the product of two (finite) totally ordered sets. Let $\mathfrak{J}$ be the set of connected posets which do not contain any chains of length $\geq 3$ and can be embedded into $\mathcal{P}$. If one computes the “sink-source compression factor” of [AENY23b] with respect to every interval subset of $\mathcal{P}$, the result is precisely $\mathbf{mult}_{\mathcal{F}}$ for =$\mathcal{F}={\{(\mathcal{X}_{i},\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}})\mid\mathcal{X}_{i}\in\mathfrak{J}\}}$. A similar result likewise holds for the “corner-complete compression factors”, while the “total compression factors” coincide with the generalized rank invariant.
@@
Let $\mathcal{P}$ be a finite poset and :=$\mathcal{F}:=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ be a family of $\mathcal{P}$-order embeddings. We define an equivalence relation on $\mathrm{E}=\bigsqcup_{i}\mathrm{E}_{i}$ by $f\sim g$ if $\overline{\operatorname{\mathrm{Im}}f}=\overline{\operatorname{\mathrm{Im}}g}$. Then the invariant $\mathbf{mult}_{\mathcal{F}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}}$ induces an isomorphism of abelian groups
@@
The $\mathcal{X}_{j}$-modules $\mathbb{I}_{\mathcal{X}_{j}}$ and $g^{*}\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}=\mathbb{I}_{g^{-1}(\overline{\operatorname{\mathrm{Im}}f})}$ are both interval modules hence indecomposable. Therefore we obtain $\left(\mathbf{mult}_{\mathcal{F}}(\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}})\right)_{g}$ is 00 or $1$ and
$\begin{array}[]{rcl}\left(\mathbf{mult}_{\mathcal{F}}(\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}})\right)_{g}=1&\Leftrightarrow&\mathcal{X}_{j}=g^{-1}(\overline{\operatorname{\mathrm{Im}}f})\\ &\Leftrightarrow&\mathcal{X}_{j}\subseteq g^{-1}(\overline{\operatorname{\mathrm{Im}}f})\\ &\Leftrightarrow&\operatorname{\mathrm{Im}}g\subseteq\overline{\operatorname{\mathrm{Im}}f}\\ &\Leftrightarrow&\overline{\operatorname{\mathrm{Im}}g}\subseteq\overline{\operatorname{\mathrm{Im}}f}.\end{array}$
@@
Generalizing the phenomenon observed for the rank invariant and sink-source compression factors, we have the following.
@@
Let $\mathcal{P}$ be a finite poset, and =$\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}$ be a family of $\mathcal{P}$-order embeddings such that there is no chain of length $\geq 3$ in $\mathcal{X}_{i}$ for each $i$. Denote $\mathrm{E}=\bigsqcup_{i}\mathrm{E}_{i}$. Then the family
$\mathcal{R}_{\mathrm{E}}:=\{\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}\mid f\in\mathrm{E}\}\subseteq\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$
forms a basis of $\mathbf{mult}_{\mathcal{F}}$.
@@ 5.3. Comparing $\mathbf{mult}$-invariants
Let $\mathcal{P}$ be a finite poset, $\mathcal{X}$ be a representation-finite poset and $\mathrm{E}$ be a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Let $\mathcal{F}_{\mathcal{P}}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ be a family of $\mathcal{P}$-order embeddings. The aim is here to compare the invariants $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$ and $\mathbf{mult}_{\mathcal{F}_{\mathcal{P}}}$. We denote by $\mathcal{F}_{\mathcal{X}}=\{(\mathcal{X}_{i},\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{X}})\}_{i}$ the family of $\mathcal{X}$-order embeddings of the posets $\mathcal{X}_{i}$. The next proposition says that it is enough to compare the invariant $\mathbf{mult}_{\mathcal{X}}$ with the invariant $\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}$.
@@
where $\beta_{(h,f,U)}=\alpha_{(g,U)}$ as soon as $h=f\circ g$. Since $f$ is an embedding the value of $\beta_{(h,f,U)}$ is well-defined.

### v3

@@ 5.1. Interval modules and the intermediate extension functor
    Now we consider $\mathcal{X}=\mathcal{X}^{\prime}_{3}$ or $\mathcal{X}^{\prime\prime}_{3}$ as follows
    $\mathcal{X}^{\prime}_{3}=$$\mathcal{X}^{\prime\prime}_{3}=112233$
    For embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$, the intermediate extension functor yields interval modules that can be viewed as replacements for the more complicated modules obtained via induction or coinduction (see Example 4.10). In particular, this provides a systematic way to avoid non-interval modules and obtain simpler building blocks for the invariant.
    We recall the description of the indecomposable $\mathcal{X}_{3}^{\prime}$- and $\mathcal{X}_{3}^{\prime\prime}$-modules from the proof of Proposition 3.12and Example 4.10: the modules $\Theta_{f}(U)$ are either simples, rectangles, or interval modules of the form ${\mathbb{I}}_{[f(1),f(2)]\cup[f(1),f(3)]}$ when $\mathcal{X}=\mathcal{X}^{\prime}_{3}$, and ${\mathbb{I}}_{[f(1),f(3)]\cup[f(2),f(3)]}$ when $\mathcal{X}=\mathcal{X}^{\prime\prime}_{3}$. In the case of a 2D-grid poset, these interval modules look as follows:
@@ 5.2. Families of embeddings
Let $\mathcal{P}$ be a poset. A family of $\mathcal{P}$-order embeddings is a finite set $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ with each $\mathcal{X}_{i}$ a finite poset (not necessarily representation finite) and each $\mathrm{E}_{i}$ a subset of $\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}}$. For $\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ a family of $\mathcal{P}$-order embeddings, we denote by $\mathrm{E}\mathrel{\mathop{\ordinarycolon}}=\bigsqcup_{i}\mathrm{E}_{i}$ and we define an invariant $\mathbf{mult}_{\mathcal{F}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}}$ by
$\left(\mathbf{mult}_{\mathcal{F}}(M)\right)_{f}\mathrel{\mathop{\ordinarycolon}}={\rm mult}(\mathbb{I}_{\mathcal{X}_{i}},f^{*}(M))\quad\textrm{for }f\in\mathrm{E}_{i}.$
@@
    Let $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=\{(\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}})\}$ for $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as in Example 3.7. We show in Example 5.13 that $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}\sqcup\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ is in natural bijection with $\{\leq_{\mathcal{P}}\}$ and via this bijection we have $\mathbf{mult}_{\mathcal{F}}=\operatorname{\mathbf{rk}}_{\mathcal{P}}$.
@@
      Let $\mathfrak{I}$ be the set of posets which embed into $\mathcal{P}$ as intervals, and for each $\mathcal{X}_{i}\in\mathfrak{I}$ let $\mathrm{E}_{i}\subseteq\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}}$ be the set of embeddings with interval image. Then, as a result of [CL18, Lemma 3.1], the generalized rank invariant of [KM21] is equal to $\mathbf{mult}_{\mathcal{F}}$ for $\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\mid\mathcal{X}_{i}\in\mathfrak{I}\}$.
@@
      Suppose that $\mathcal{P}$ is the product of two (finite) totally ordered sets. Let $\mathfrak{J}$ be the set of connected posets which do not contain any chains of length $\geq 3$ and can be embedded into $\mathcal{P}$. If one computes the “sink-source compression factor” of [AENY23b] with respect to every interval subset of $\mathcal{P}$, the result is precisely $\mathbf{mult}_{\mathcal{F}}$ for $\mathcal{F}={\{(\mathcal{X}_{i},\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}})\mid\mathcal{X}_{i}\in\mathfrak{J}\}}$. A similar result likewise holds for the “corner-complete compression factors”, while the “total compression factors” coincide with the generalized rank invariant.
@@
Let $\mathcal{P}$ be a finite poset and $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}_{i}$ be a family of $\mathcal{P}$-order embeddings. We define an equivalence relation on $\mathrm{E}=\bigsqcup_{i}\mathrm{E}_{i}$ by $f\sim g$ if $\overline{\operatorname{\mathrm{Im}}f}=\overline{\operatorname{\mathrm{Im}}g}$. Then the invariant $\mathbf{mult}_{\mathcal{F}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}}$ induces an isomorphism of abelian groups
@@
The $\mathcal{X}_{j}$-modules $\mathbb{I}_{\mathcal{X}_{j}}$ and $g^{*}\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}=\mathbb{I}_{g^{-1}(\overline{\operatorname{\mathrm{Im}}f})}$ are both interval modules hence indecomposable. Therefore we obtain $\left(\mathbf{mult}_{\mathcal{F}}(\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}})\right)_{g}$ is 0 or $1$ and
$\begin{array}[]{rcl}\left(\mathbf{mult}_{\mathcal{F}}(\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}})\right)_{g}=1&\Leftrightarrow&\mathbb{I}_{\mathcal{X}_{j}}=\mathbb{I}_{g^{-1}(\overline{\operatorname{\mathrm{Im}}f})}\\ &\Leftrightarrow&\mathcal{X}_{j}=g^{-1}(\overline{\operatorname{\mathrm{Im}}f})\\ &\Leftrightarrow&\mathcal{X}_{j}\subseteq g^{-1}(\overline{\operatorname{\mathrm{Im}}f})\\ &\Leftrightarrow&\operatorname{\mathrm{Im}}g\subseteq\overline{\operatorname{\mathrm{Im}}f}\\ &\Leftrightarrow&\overline{\operatorname{\mathrm{Im}}g}\subseteq\overline{\operatorname{\mathrm{Im}}f}.\end{array}$
@@
We now show that the perspective of iterated embeddings, combined with the intermediate extension functor, yields explicit bases for the associated invariants. In particular, these bases consist of interval modules and generalize the classical rectangle-module basis for the rank invariant.
@@
Let $\mathcal{P}$ be a finite poset, and let $\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})\}$ be a family of $\mathcal{P}$-order embeddings such that there is no chain of length $\geq 3$ in $\mathcal{X}_{i}$ for each $i$. Denote $\mathrm{E}=\bigsqcup_{i}\mathrm{E}_{i}$. Then the invariant $\mathbf{mult}_{\mathcal{F}}$ admits a basis
$\mathcal{R}_{\mathrm{E}}\mathrel{\mathop{\ordinarycolon}}=\{\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}\mid f\in\mathrm{E}\}\subseteq\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$
given by the modules obtained via the intermediate extension functor from the characteristic indecomposable modules of the $\mathcal{X}_{i}$.
@@ 5.3. Comparing $\mathbf{mult}$-invariants
Let $\mathcal{P}$ be a finite poset, $\mathcal{X}$ be a representation-finite poset and $\mathrm{E}$ be a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. As mentionned before, in general the rank of the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$ is much smaller than $|\mathrm{E}\times\operatorname{\mathrm{Ind}}\mathcal{X}|$. The aim of this section is to find a family $\mathcal{F}=\{(\mathcal{X}_{i},\mathrm{E}_{i})_{i}\}$ of $\mathcal{P}$-order embeddings with $\sum_{i}|\mathrm{E}_{i}|\leq|\mathrm{E}\times\operatorname{\mathrm{Ind}}\mathcal{X}|$ and with $\mathbf{mult}_{\mathcal{F}}\sim\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$. We start with the example coming from 3.9 to illustrate the strategy.
Example 5.13.
Let $\mathcal{P}$ be a finite poset and $\mathcal{F}=\{(\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}})\}$ given as follows.
$\mathcal{X}_{1}=$$\mathcal{X}_{2}=11$22
One can define a bijection $\{\leq_{\mathcal{P}}\}\to\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}\sqcup\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ sending $(a,a)\in\{\leq_{\mathcal{P}}\}$ to $f_{a}\in\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}$ with $f_{a}(1)=a$ and $(a,b)$ to $f_{ab}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}_{2}\to\mathcal{P}$ with $f_{ab}(1)=a$ and $f_{ab}(2)=b$. Then it is straightforward to check that we have a commutative square
${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{F}}}$${\mathbb{Z}^{\mathrm{E}_{1}\sqcup\mathrm{E}_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{\mathbf{rk}}_{\mathcal{P}}}$${\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}}$
showing that $\mathbf{mult}_{\mathcal{F}}\sim\operatorname{\mathbf{rk}}_{\mathcal{P}}$. Combining this with Corollary 3.8, we obtain that if $\mathcal{P}$ is connected with at least $2$ elements then we have an equivalence
$\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}\sim\mathbf{mult}_{\mathcal{F}}.$
As noted in Remark 3.9, computing $\mathbf{mult}_{\mathcal{F}}$ is roughly three times more efficient than computing $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}.$
This equivalence can also be seen directly using the following arguments. Consider the family of $\mathcal{X}_{2}$-order embeddings
$\mathcal{F}_{\mathcal{X}_{2}}\mathrel{\mathop{\ordinarycolon}}=((\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}_{2}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}_{2}})).$
Then we have $\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}_{2}}=\{f_{1},f_{2}\}$ with $\operatorname{\mathrm{Im}}f_{x}=\{x\}$, and $\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}_{2}}=\{{\rm id}\}$. We immediately check for $M$ in $\operatorname{\mathrm{mod}}\mathcal{X}_{2}$
$\left(\mathbf{mult}_{\mathcal{F}_{\mathcal{X}_{2}}}(M)\right)_{f}=\left\{\begin{array}[]{cl}\dim M_{x}&\textrm{ if }f=f_{x}\\ \operatorname{\mathrm{rk}}M_{1\to 2}&\textrm{ if }f={\rm id}\end{array}\right$.
Then one has a commutative square
${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X}_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{X}}}$${\mathbb{Z}^{\{\mathbb{I}_{\{2\}},\mathbb{I}_{\{1,2\}},\mathbb{I}_{\{1\}}\}}}$${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X}_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}}$${\mathbb{Z}^{\{f_{1},f_{2},{\rm id}\}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi=\begin{pmatrix}0&1&-1\\ 0&0&\;\;\;1\\ 1&0&-1\end{pmatrix}}$
Now fix a $g\in\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}$. Since $\mathcal{P}$ is connected and has at least $2$ elements, this embedding $g$ can be written as $g=f\circ f_{x}$ with $f\in\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ and $f_{x}\in\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}_{2}}$ ($x=1,2$). Then one can write
$(\mathbf{mult}_{\mathcal{F}}(M))_{g}$ $=(\mathbf{mult}_{\mathcal{X}_{1}}(g^{*}M))_{\mathbb{I}_{\{1\}}}$
$=(\mathbf{mult}_{\mathcal{X}_{1}}((f\circ f_{x})^{*}M))_{\mathbb{I}_{\{1\}}}$
$=(\mathbf{mult}_{\mathcal{X}_{1}}(f_{x}^{*}(f^{*}M)))_{\mathbb{I}_{\{1\}}}$
$=(\mathbf{mult}_{\mathcal{F}_{\mathcal{X}_{2}}}(f^{*}M))_{f_{x}}$
$=\left(\Phi^{-1}\circ\mathbf{mult}_{\mathcal{X}}(f^{*}M)\right)_{f_{x}}$
$=\left(\Phi^{-1}\circ\mathbf{mult}_{\mathcal{X},\mathcal{P}}(M)_{(f,-)}\right)_{f_{x}}.$
A similar argument can be used if $g\in\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ which gives a direct proof that the invariants $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ and $\mathbf{mult}_{\mathcal{F}}$ are equivalent.
Using the same kind of arguments, one can show the following more general result.
@@
Let $\mathcal{P}$ be a finite poset, $\mathcal{X}$ be a representation-finite poset and $\mathrm{E}$ be a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Let $\mathcal{F}_{\mathcal{P}}=\{(\mathcal{X}_{i},\mathrm{E}_{i})_{i}\}$ be a family of $\mathcal{P}$-order embeddings. We denote by $\mathcal{F}_{\mathcal{X}}=\{(\mathcal{X}_{i},\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{X}})\}_{i}$ the family of $\mathcal{X}$-order embeddings of the posets $\mathcal{X}_{i}$.
@@
where $\beta_{(h,f,U)}=\alpha_{(g,U)}$when $h=f\circ g$. Since $f$ is an embedding the value of $\beta_{(h,f,U)}$ is well-defined.